kaylashultz:

The equation d =(m/d) can be used to calculate the density, d, of an object with mass, m, and volume, V. Which is an equivalent equation solved for V?

kaylashultz:

answer choices: dm = V, (d/m)= V, (m/d)=V, and m-d=V

jhonyy9:

Density is defined as mass per unit volume. d = m/V so from this result : V = m/d hope helped